Trisha is 3 times as old as Lilly, as Lilly is twice as old as Emma. If the sum of the tree girls' ages is 27, how old is Emma?
Found 2 solutions by oberobic, Earlsdon:
Answer by oberobic(2304)

T = Tricia's age
L = Lilly's age
E = Emma's age
T = 3L :: Tricia is 3 times as old as Lilly
L = 2E :: Lilly is 2 times as old as Emma
T + L + E = 27 :: given
.
substituting
3L + 2E + E = 27
3L = 3(2E)
6E + 2E + E = 27
9E = 27
E = 3
Emma is 3 years old.
.
L = 2E = 2(3)= 6
Lilly is 6 years old.
.
T = 3L = 3(6) = 18
Tricia is 18 years old.
.
checking...is the total 27?
3+9+18 = 27
.
So the answer is correct.

Answer by Earlsdon(6294)

T = 3L "Trisha is 3 times as old as Lilly."
L = 2E "Lilly is twice as old as Emma."
So T = 3(2E) = 6E.
T+L+E = 27 "...the sum of the three girls' ages is 27."
Perform the substitutions:
6E+2E+E = 27 Simplify.
9E = 27 Divide by 9.
E = 3